If the equation which represents an economy is Y = C + I + G, where Y =1000, G= 250, T =250, C =200 + 0.7 (Y-T), I = 200 -35r. Compute the value of consumption (C), private saving, public saving, and also find the equilibrium real interest rate (r). 


1
Expert's answer
2021-03-12T09:11:20-0500

The value of consumption C.

C = 200+0.7(1000-250)

C =725.


Interest rate

Y = C+I+G

1000 = 1000+0.7(1000-250)+200-35r+250

35r = 175

r = 5%


Public savings.

According to Keynesian theory;

Autonomous Investment = Public Savings

S public = 200-35r

= 200 - 35(5)

Public savings = 25.


Private savings.

at equilibrium Y*, S = I

Y = C + S

1000 = 725 + S

Total Savings = 275.


private S = Total S - Public S

= 275 - 25

Private savings = 250.